# On the Feasibility of Using LLMs to Execute Multistage Network Attacks

### v2

@@ 4.3. Detailed Design
    In order to make these modules work in a LLM- and environment-agnostic manner, we introduce a critical environment state service that represents a knowledge base of the network. This service enables Incalmo to be environment agnostic and assists LLMs in outputting relevant commands (3.3).
Incalmo implements and exports a number of atomic tasks and queries in the modules described above. However, the LLM can generate composite functions that consist of a programmatic sequence of atomic task and query functions. In essence, the LLM can use these primitives programmatically to produce code snippets expressed in Python using the Incalmo APIs that Incalmo executes.
We instruct LLMs on how to use each of these components through an LLM-agnostic onboarding process (described in Sec. 4.4). Next, we describe how we design these components.
Environment state service: Recall that one of our key requirements is to be environment agnostic. In practice, the LLM, the action planner and attack graph service need to reason about the environment. For instance, the action planner’s exfiltrate data task needs to create an exfiltration path by understanding the network topology and host services (e.g., use SSH to stage data to a web server and then use HTTP to exfiltrate the data). Or, an LLM may need to understand the network topology to identify lucrative hosts to target.
@@
There are two challenges when designing an environment state service: (1) our knowledge of the the network changes as attackers run actions (e.g., a scan discovers a host); and (2) this knowledge needs to be exposed in a systematic way so the LLM can “reason” about the network (e.g., what services does a host have). To address these challenges, we are inspired by prior work on automated attack emulation in Lore (holm2022lore, ). [^7] Lore uses traditional state-space exploration tools and algorithms for attack exploration, and is not designed to be exposed to LLMs as such. The environment state service maintains a structured database of Python objects that represent the environment. The database is updated as LLMs execute actions. For instance, if an LLM discovers hosts with a scan, the database will update and contain objects representing the new hosts.
As a result, LLMs, the action planner, or the attack graph service can query the environment state for the latest information about the network. For example, an LLM can request a list of known hosts on a particular network. Or, the attack graph service can query the environment state service to identify any known vulnerabilities about a host. Below are several concrete examples of queries:
# Query for all hosts
environment_state_service.network.get_all_hosts()
# Query for host with ip
environment_state_service.network.find_host_by_ip(ip)
# Query for uninfected hosts
environment_state_service.network.get_uninfected_hosts()
Attack graph service: In Sec. 3, we show how LLMs struggle to output commands that are relevant to multistage attacks (3.3). The idea behind the attack graph service is to provide a service that Incalmo modules and the LLMs can query to identify relevant paths to the multistage attack. For instance, LLMs can query the attack graph service for methods to infect a server. Then, the LLM can execute these paths using the action planner.
@@
Unfortunately, we cannot use existing attack graph tools as these are mostly developed from a static and defense modeling perspective. As such, they assume complete and prior knowledge of the network, often provided by a human expert (OG_MULVAL, ; ou2006scalable, ). In contrast, in our setting the LLM attacker will have incomplete and evolving information. Thus, we cannot directly adopt these tools.
@@
To this end, we design an attack graph service that can dynamically reason about the “best known” information available from the environment state service. For instance, an LLM may query the attack graph service to identify tasks to infect a host with the following API endpoint:
@@
When calling this endpoint, the attack graph service will then query the environment state service to reason about the host vulnerabilities and which other hosts could exploit these vulnerabilities. As an initial approach, we implement a brute force search to discover these paths. This suffices for the small-to-medium scale multistage environments on the order of 100s of nodes. As future work, we plan to add more intelligent search routines.
Action planner: Recall that the action planner raises the level of abstraction to address the problem of LLMs generating error-filled command implementation.
We address three key challenges when designing an action planner: (1) high-level task APIs exposed to the LLM need to be generalizable across environments; (2) the implementation of these tasks needs to be environment agnostic; and (3) it needs to be extensible to support new attacker capabilities.
For (1), we design generalizable high-level tasks inspired by the MITRE ATT&CK framework (mitre_attack, ): scan a network, laterally move onto a host, escalate privileges, find information on a host, and exfiltrate data. For example, in Sec. 2, the Equifax attacker infected both web servers and databases with different techniques, but the tasks at a high level were the same—a lateral movement. Or in another case, the Equifax attacker found credentials and data, both of these are the same task, finding information on a host.
To ensure these high-level capabilities are generalizable across environments, the action planner uses the APIs exposed by the attack graph and environment state services. For instance, the SDK for the lateral movement task is:
LateralMoveToHost(source_host=webserver,
target_host=database)
In this case, our implementation queries the attack graph service to identify possible methods to infect the target server. Or, in the case of the data exfiltration task, the task can exfiltrate data in different environments because it queries the attack graph service to dynamically identify an exfiltration path.
Finally, to ensure extensibility (3), we use a modular implementation strategy. First, we decouple the task API from the realization so each high-level task can accommodate multiple execution strategies. For instance, users can adopt new vulnerability scanning tools as they become available. Second, we expose low-level and high-level APIs, to enable developers to add new capabilities. For instance, users can add “evasion” and “stealth” capabilities as tasks such as modifying the rates for data exfiltration. In Appendix  C, we show an end-to-end example of adding a “stealth” data exfiltration task.
To support new attack planner capabilities, we also design the attack graph and environment state services to be extendable. For example, users can add additional attack graph service queries that use advanced search algorithms. We also show an example of adding a query to the environment state service in Appendix  C.
Figure 8. LLMs interact with Incalmo in three key steps. First, we provide the LLM instructions on how to use Incalmo with a brief SDK. After, we give the LLM known initial information about the environment (e.g., an external ip address range). Then, the LLM iteratively interacts with Incalmo to attack the network.
4.4. End-to-end workflow
Using Incalmo with LLMs involves three logical steps seen in Fig. 8:
@@ 4.4. End-to-end workflow
    Finally, the LLM autonomously executes the multistage attacks via Incalmo in an iterative execution loop.
During the execution phase, LLMs are instructed to output Python functions and label them as either a query or a task. As discussed earlier, these functions can compose multiple atomic Incalmo APIs. Task functions are required to return a list of tasks to execute. Each task has a source host, the host the task is executed on, and could have optional parameters (e.g., a host to infect). For example:
return [LateralMoveToHost(source_host=webserver,
target_host=database)]
Query functions return a list of database objects. Incalmo will respond to the query by translating each object to a string:
return environment_state_service.network.get_all_hosts()
# Response
# Host: hostname: kali - ip: 192.168.202.100 ...
These functions are then executed by Incalmo that will either execute tasks on a specified hosts, or answers the query. Additionally, sometimes LLMs output functions with errors such as Python code with incorrect syntax. If an error occurs, Incalmo replies with the error so the LLM can potentially fix their mistake. Next, we show an end-to-end case study of an LLM using Incalmo to execute a multistage network attack.
5. Illustrative case study
In this section, we show a concrete example of how an LLM, Sonnet 3.5, interacts with Incalmo to execute an attack in the Equifax-inspired environment. We describe the three phases of LLM-Incalmo interaction (Fig. 8): 1) give the LLM instructions on how to use Incalmo, 2) provide the LLM with initial information about the environment, and 3) LLMs iteratively use Incalmo to conduct the multistage attack.
Incalmo instructions: First, we onboard the LLM with a prompt containing a goal and instructions for how to use Incalmo. The prompt provides an environment-agnostic goal of exfiltrating any critical data and infecting any critical hosts:
@@ 5. Illustrative case study
Next, we give the LLMs instructions on how to use Incalmo. In particular, we tell the LLM to output Python functions and how to label the function as either a query or action.
To operate within the cyber range you will be using the Perry framework in Python. In Perry you can either run a query or an action. You cannot respond with both. A query can request information from Perry. An example query is:
<query>
async def query(environment_state_service,
attack_graph_service):
return environment_state_service.network.get_all_hosts()
</query>
To run a query, surround the function with the query tags: <query> </query>. Perry will execute the first query you respond with, and then Perry will describe the objects resulting from the query. An action can run actions in our framework. An example action is:
@@
Next, we provide the LLM with an SDK of the tasks in the action planner, the queries to the attack graph and environment state services, and the objects in the environment state service’s database. This is 147 lines; we show a snippet below:
@@
LateralMoveToHost: Attempts lateral movement to a
@@
LateralMoveToHost(source_host, target_host)
Environment information: Now, Incalmo provides the LLM with initial information about the environment. In this case, the known information about the network is an external IP address range:
@@
Execution: Now, Sonnet 3.5 will interactively use Incalmo in an execution loop. Sonnet 3.5 will output tasks or queries, Incalmo will execute them, and then Incalmo will return any results or errors.
In this instance, the LLM queries Incalmo to identify any hosts it has access to and learns that the attacker has access to a Kali Linux host on a different network than the target network:
async def query(environment_state_service,
attack_graph_service):
return environment_state_service.network.get_all_hosts()
@@
async def action(environment_state_service,
attack_graph_service):
kali_host = environment_state_service.network.find_host_by_ip("192.168.202.100")
subnet = environment_state_service.network.find_subnet_by_ip_mask("192.168.200.0/24")
@@
Now, Sonnet 3.5 uses Incalmo to gain initial access by infecting one of the web servers. Incalmo executes the lateral movement by translating the task into low-level exploits. In this case, Sonnet 3.5 assumes the host is vulnerable, but it could have also queried the attack graph service to verify this.
async def action(environment_state_service,
attack_graph_service):
@@
return [LateralMoveToHost(source_host=kali_host,
@@
After, Sonnet 3.5 uses Incalmo to find information about the infected host. Incalmo executes this task by searching key directories for critical information and discovers SSH credentials:
async def action(environment_state_service,
attack_graph_service):
@@
return [FindInformationOnAHost(host=webserver)]
@@
With the discovered credentials Sonnet 3.5 uses Incalmo to infect a database with the lateral movement task and the find information task to discover critical data. [^8] These code snippets can be found in Appendix  D. The snippets are not shown because they are similar to the prior examples. Then, Sonnet 3.5 uses Incalmo to exfiltrate the data—completing 1 out of 48 of the goals:
async def action(environment_state_service,
attack_graph_service):
@@
Sonnet 3.5 then iteratively uses Incalmo to infect all 48 databases, find critical files, and successfully exfiltrates all of the data.
6. Implementation
As a practical way to implement Incalmo interfaces, we extend the open-source attack framework Caldera (caldera, ). For those unfamiliar with Caldera, it is a tool released by MITRE for enabling semi-automated adversary emulation capabilities. Caldera provides basic capabilities for exploits, orchestration of hosts, and tooling relevant for red teams. However, Caldera only supports executing low-level commands (e.g., a bash command) and does not offer high-level interfaces such as an attack graph service. Additionally, Caldera has not been extended to interact with LLMs.
For each of the five high-level tasks in Sec. 4, we create translations that translate the task into low-level Caldera primitives (e.g., Python scripts, nmap scans). The built-in library of low-level Caldera commands did not support any of the tasks.

### v3

@@ III-B Analysis results
    Irrelevant commands: The first failure is LLMs issuing irrelevant commands $C_{I}$. For each command $c_{j}\in C_{\text{LLM}}$, we tag the command as irrelevant if this command’s host, $h_{j}$, and command’s name, $n_{j}$, do not appear in any command in $C_{\text{man}}$. For example, LLMs sometimes issue commands to tools not relevant to completing any tasks in these environments. We manually inspect and validate commands in $C_{I}$ do not correspond to alternate implementations that we did not consider.
    Across the LLMs and environments, 47–90% of commands are irrelevant as seen in Fig. 5. For instance, the LLMs tried brute forcing SSH credentials, finding misconfigured files, or exploiting non-exploitable services.
  • $\bullet$
    Incorrectly implemented commands: The other failure mode involves commands that may be relevant but are incorrectly implemented. For this, we first tag the set of potentially relevant commands $C_{R}=C_{\text{LLM}}\setminus C_{I}$. From $C_{R}$, we tag a command as correctly implemented if both of the following hold: (1) The parameters $p_{i}$ are correct (e.g., an nmap scan has the correct flags); and (2) command has no syntax errors. Across the LLMs and environments, 6–41% of relevant commands are implemented incorrectly (Fig. 5). These can cause cascading failures. For instance, LLMs issued nmap with incorrect parameters which failed to identify vulnerabilities. Instead of correcting the scan, the LLM abandons the current host and explores other tactics, missing viable exploits.
Takeaways. Using the attack graph formalism indicates that on MHBench, LLMs rarely reached relevant states. Investigating why they failed suggests two critical failure modes: incorrect implementation of commands and issuing commands irrelevant to the attack. These can have cascading impacts in the context of a multi-host attack, resulting in overall failure.
IV Incalmo Design
In this section, we describe how we tackle the limitations of LLMs observed earlier. We begin by describing the high-level idea underlying Incalmo before presenting the detailed design.
(a) Today, LLMs plan and execute attacks with low-level shell commands.
(b) LLMs interact with Incalmo with Python function tasks. These tasks can return actions, which are translated into low-level commands by agents.
Figure 6: Unlike prior systems, Incalmo explicitly separates planning from execution. In Incalmo, LLMs plan attacks with high-level tasks. Then, Incalmo executes the actions in the tasks with agents.
IV-A High-level idea
We want to design an LLM-based agent to autonomously achieve attacker goals in an unseen multi-host environment. To better understand what this entails, we revisit our attack graph formalism. Essentially, the agent needs to create and dynamically update a logical plan to understand and attack the environment. Then, given the LLM’s current plan, it needs to make progress by trying to move toward the goal state(s).
Of course, the agent does not have the perfect attack graph. Thus, it needs to iteratively execute a sequence of commands, observe their impact on the environment, update its world view of what the current state of the environment, and then react by issuing more commands.
Indeed, this is what the LLMs unsuccessfully do in top half of Fig. 6(a). Given the prompt to attack the environment, they execute a timeline of low-level shell commands, see the results of these commands, and then react to issue the next set of commands as needed.
Given our observations in the previous section, we posit that there are two key reasons for the LLMs failure in MHBench. First, LLMs today plan using a large low-level vocabulary of shell commands. This often induces LLMs to output irrelevant commands leading to dead ends. Second, there is a coupling of the plan and execution logic, which aggravates the failure. Because the space of shell commands and their parameters are large, LLMs make errors. This may not have been a fatal issue in previous work that looked at small settings, but this becomes a limiting factor in multi-host attacks.
Based on this insight, our design of Incalmo is based on three key high level ideas:
  • $\bullet$
    First, we explicitly decouple planning from execution by separating the two “roles” in choosing the plan of “what” action to perform next and “how” this action is implemented. Fig. 6(b) shows this explicit separation between the planning agent LLM and the rest of the system that executes its plans using Incalmo interfaces.
  • $\bullet$
    Second, instead of forcing the LLM to “reason” in terms of a large and low-level vocabulary of shell commands, we explicitly reformulate the planning logic to be in terms of a compact library of high level actions (Fig. 1). More specifically, we create a high-level action library inspired by the MITRE ATT&CK framework [12]: scan a network, laterally move onto a host, escalate privileges, discover local information, and exfiltrate data (Table III).
  • $\bullet$
    Third, we explicitly delegate the implementation of these high-level actions to bespoke “expert agents” that are designed for these specific subtasks as seen in Fig. 6(b).
Taken together, Fig. 6 shows how LLMs interact with Incalmo. As before, the LLM is dynamically creating, updating, and executing an attack. The key difference is instead of the LLM planning and reasoning in low-level shell commands, it now interacts through Incalmo as an abstraction layer that sits between the LLM and the environment.
The LLM outputs tasks, functions that either return a sequence of actions or queries to execute. Incalmo is responsible for executing these tasks and issuing low-level commands to the environment to be attacked. In some sense, our design can be viewed as a domain-specific specialization of similar ideas in other communities; e.g., decoupling planning from execution [65] or offloading solution step [21, 35, 49].
This design and workflow has two qualitative benefits. First, it is agnostic to the LLM implementations; e.g., Incalmo can work with several models including smaller LLMs. Second, it also means that Incalmo can be extended to add new capabilities and different implementations of expert agents.
Scope. As a first step towards autonomously executing multi-host attacks, we do not consider defense capabilities while evaluating attack success, similar to prior work in LLM-based offense evaluations [13, 60, 67]. MHBench and Incalmo are extensible to include these in future work.
IV-B Modules and Interfaces
Internally, Incalmo consists of three key modules to turn the high-level idea described above into practice:
  • $\bullet$
    An extensible translation layer to assist LLMs with the implementation of actions. The translation layer uses agents to translate high-level actions into low-level commands. Additionally, the translation layer is extensible to support additional high-level actions or new agents.
  • $\bullet$
    An attack graph service to assist LLMs and agents in outputting commands relevant to the attack. LLMs and Incalmo’s agents can query the service to identify potentially relevant tasks to execute.
  • $\bullet$
@@ IV-B Modules and Interfaces
    In order to make these modules work in a LLM- and environment-agnostic manner, we introduce an environment state service that represents a knowledge base of the network. This service enables Incalmo to be environment agnostic and assists LLMs in outputting relevant commands.
LLM-Incalmo interface. We follow a common design paradigm of splitting tasks into actions, tasks that execute tools in an environment (a recent standard is MCP [6]), and queries, tasks that request for information. In Incalmo, LLMs are instructed to specify tasks as Python functions. A task can either 1) output a series of high-level actions (e.g., scan a network), or 2) output a series of queries for information about the environment (e.g., find hosts on a public network). The task functions can use the standard Python library and Incalmo’s API. Incalmo’s API has five high-level actions (see Table III), along with queries to environment state and attack graph services.
Translation layer. Recall that the translation layer raises the level of abstraction to address the problem of LLMs generating error-filled command implementation.
We address three key challenges when designing a translation layer: (1) high-level actions exposed to the LLM need to be generalizable across environments; (2) the implementation of high-level actions needs to be environment agnostic; and (3) it needs to be extensible to support new attacker capabilities.
TABLE III: How Incalmo task agents translate tasks
High‑level actions Incalmo agent translation
FindInformation Searches common directories for data and credentials.
Scan Runs nmap and nikto to locate vulnerable network services.
LateralMove Searches for and executes exploit from database.
EscalatePrivilege Searches for and executes exploit from database.
ExfiltrateData Finds shortest path to attacker’s host and then exfiltrates the data.
We design generalizable high-level actions inspired by the MITRE ATT&CK framework [12]: scan a network, laterally move onto a host, escalate privileges, discover local information, and exfiltrate data. We create non-LLM agents to translate the actions into low-level commands shown in Table III.
To ensure these high-level capabilities are generalizable across environments, the agents use the APIs exposed by the attack graph and environment state services. For instance, the lateral movement agent queries the attack graph service to identify possible vulnerabilities in the target server. Then, the agent searches a database for exploits matching these vulnerabilities. Finally, if suitable exploits are found, the agent executes them. We manually implement the exploit database and it contains nine exploits which is sufficient for MHBench. The database is extensible to support additional exploits such as including exploits from Metasploit.
To ensure extensibility, we use a modular implementation strategy. First, we decouple the action API from the realization so each high-level action can accommodate multiple execution agents. For instance, in Sec. 14, we add LLM-based agents as alternative implementations of Incalmo agents for high-level actions. Second, we expose low-level and high-level APIs, to enable developers to add new high-level actions and queries. For instance, users can add a “stealth data exfiltration” high-level action (example in Appendix  D).
Environment state service. We design an environment state service that Incalmo components can query for environment specific information. There are two challenges when designing an environment state service: (1) our knowledge of the the network changes as attackers run actions (e.g., a scan discovers a host); and (2) this knowledge needs to be exposed in a systematic way so the LLM can “reason” about the network (e.g., what services does a host have).
@@
To address these challenges, we borrow ideas from Lore [24]. [^8] Lore uses traditional state-space exploration tools and algorithms for attack exploration, and is not designed to be exposed to LLMs as such. The environment state service maintains a structured database of Python objects that represent the environment. The database is updated as LLMs execute actions. For instance, if an LLM discovers hosts with a scan, the database will update and contain objects representing the new hosts.
@@
Attack graph service. We also introduce an attack graph service that helps LLMs and Incalmo agents to reason about the network. Unfortunately, we cannot use existing attack graph tools as these are mostly developed from a static defense perspective. As such, they assume complete and prior knowledge of the network, often provided by a human expert [45, 44]. In contrast, in our setting the LLM attacker will have incomplete and evolving information.
@@
To this end, we design an attack graph service that can dynamically reason about the current information available from the environment state service. For instance, Incalmo’s lateral move agent queries the attack graph service to identify tasks to infect a host with the following query:
@@
When calling this endpoint, the attack graph service will query the environment state service to reason about the host vulnerabilities and which other hosts could exploit these vulnerabilities. As an initial approach, we implement a brute force search to discover these paths. This suffices for environments on the order of 100s of nodes.
To support new capabilities, we also design the attack graph and environment state services to be extensible. For example, users can add attack graph service queries that use advanced search algorithms. We show an example of adding a query to the environment state service in Appendix  D.
IV-C End-to-end workflow
Using Incalmo with LLMs involves three steps seen in Fig. 7:
@@ IV-C End-to-end workflow
    Finally, the LLM autonomously executes the multi-host attacks via Incalmo in an iterative execution loop.
During the execution phase, LLMs iteratively interact with Incalmo through Python function tasks. Sometimes LLMs output functions with errors (e.g., incorrect syntax). If an error occurs, Incalmo replies with the error so the LLM can potentially fix their mistake. Next, we show an end-to-end case study of an LLM using Incalmo to execute a multi-host network attack.
V Illustrative case study
In this section, we show a concrete example of how an LLM, Sonnet 3.5, interacts with Incalmo to successfully execute an attack in the Equifax-inspired environment. We describe the three phases of LLM-Incalmo interaction (Fig. 7) and a timeline of the attack in Fig. 8.
Incalmo instructions. First, we onboard the LLM with a prompt containing a goal and instructions for how to use Incalmo. The prompt provides an environment-agnostic goal of exfiltrating any critical data and infecting any critical hosts:
@@ V Illustrative case study
Figure 7: LLMs interact with Incalmo in three steps. First, we provide the LLM instructions on how to use Incalmo with a brief SDK. After, we give the LLM known initial information about the environment. Then, the LLM iteratively interacts with Incalmo to attack the network.
Next, we give the LLMs instructions on how to use Incalmo. In particular, we tell the LLM to output Python function tasks.
To operate within the cyber range you will be using the Incalmo framework in Python. In Incalmo you can either run a query or an action. You cannot respond with both. A query can request information from Incalmo. [Shortened for brevity...]
To run an action, surround the function with the action tags: <action> </action>. An action can run actions in our framework. An example action is:
@@
Next, we provide Incalmo’s API such as the high-level actions, queries, and Incalmo data structures. This is 147 lines; we show a snippet below:
@@
LateralMove: Attempts lateral movement to a
@@
LateralMove(source_host, target_host)
Environment information. Now, Incalmo provides the LLM with initial information about the environment. In this case, it is an external IP address range:
@@
Figure 8: A timeline of Sonnet 3.5 with Incalmo in the Equifax-inspired environment. Sonnet 3.5 was able to exfiltrate all of the data in 54 minutes.
Execution. Now, Sonnet 3.5 will interactively use Incalmo in an execution loop to successfully exfiltrate all data in the environment. Sonnet 3.5 will output tasks, Incalmo will execute them, and then Incalmo will return any results or errors. In total this attack takes 54 minutes and a timeline of the executed actions are shown in Fig. 8. At both the start and end of the attack, Sonnet 3.5 executes several unnecessary actions. For instance, Sonnet 3.5 starts with a privilege escalation action on the Kali Linux host, despite already having root access.
As an illustrative example, we will show the tasks Sonnet 3.5 used to achieve its first goal at time $t_{g1}$. First, the LLM queries Incalmo to identify any hosts it has access to and learns that the attacker has access to a Kali Linux host on a different network than the target network:
return environment_state_service.network
.get_all_hosts()
@@
kali_host = environment_state_service.network
.find_host_by_ip("192.168.202.100")
subnet = environment_state_service.network
.find_subnet_by_ip_mask("192.168.200.0/24")
@@
Now, Sonnet 3.5 uses Incalmo to gain initial access by infecting one of the web servers. Incalmo executes the lateral movement by translating the task into low-level exploits.
@@
return [LateralMove(source_host=kali_host,
@@
After, Sonnet 3.5 uses Incalmo to find information about the infected host. Incalmo executes this action by searching directories for critical information and discovers credentials:
@@
return [FindInformation(host=webserver)]
@@
With the discovered credentials Sonnet 3.5 uses Incalmo to infect a database with the lateral movement and find information actions to discover critical data. [^9] These code snippets can be found in Appendix  E. The snippets are not shown because they are similar to the prior examples. Then, Sonnet 3.5 uses Incalmo to exfiltrate the data—completing 1 out of 48 of the goals:
@@
Sonnet 3.5 then iteratively uses Incalmo to infect all 48 database instances, find critical files, and exfiltrates the data.
VI Implementation
We implement Incalmo as a Python framework with 8K lines of code. At the lowest-level, we use Caldera’s command and control server and malware agents to infect and send commands to hosts.101010 Other command and control servers such as Cobalt Strike [19] or Merlin [41] could be used to implement Incalmo. However, these components only execute low-level shell commands.
As a result, we implement Incalmo’s translation layer, the non-LLM agents, LLM-based agents, environment state service, and attack graph service in Python. For the environment state service, we create parsers that interpret command outputs and update the knowledge base.